OpenStudy (anonymous):

can anyone help me with this... x^2/3 + 4x^1/3 + 3 = 0 How do I solve this?

OpenStudy (anonymous):

Factorize it as : (x^(1/3)+1) (x^(1/3)+3) = 0

OpenStudy (campbell_st):

well the equation is \[\sqrt[3]{x^2} + 4\sqrt[3]{x} + 3 = 0\] use a substitution and let \[u = x^ (1/3)\] then the problem is \[u^2 + 4u + 3 = 0\] solve for u ((u+1)(u+3)=0 u = -1,-3 then find \[x = \sqrt[3]{(-1)} \] and \[x = \sqrt[3]{-3}\] are the solutions